OpenStudy (anonymous):

Write an equation for each line in slope intercept form. Slope = ½ and the y-intercept is (0,6) Rewrite the equations In slope intercept form. Graph each line. -4x + 6y =18 3y + x = 5 Write an equation for each line in point-slope form Slope =5, through (1,-1) through (2,3) and (3, 5)

OpenStudy (triciaal):

slope intercept form for the equation of the line y = m x + b where m is the slope and b is the intercept for the point (x, y)

OpenStudy (triciaal):

y-intercept is (0,6) b = value of y when x = 0 = 6 y = 1/2(x) + 6

OpenStudy (triciaal):

Rewrite the equations In slope intercept form write in the form y = remember to keep the balance of the equation

OpenStudy (triciaal):

Graph each line. you only need 2 points to graph a line. the easiest to use are when y = 0 and when x = 0

OpenStudy (triciaal):

point-slope form y - y1 = m ( x - x1) where m is the slope and the points are (x1,y1) and (x, y)

OpenStudy (anonymous):

y-6 =1/2(x -0) y=1/2x -6.5 would this be right?

OpenStudy (anonymous):

y=1/2x + 6

OpenStudy (triciaal):

as posted above

OpenStudy (anonymous):

Oh okay, thanks a lot!

OpenStudy (triciaal):

read through all the post and let me know if you need more. do you want to post a sketch of the graph etc

OpenStudy (anonymous):

That would be helpful, if you don't mind. : D